Which value of r gives the line passing through (3,2) and (r,-4) a slop of 3?

Answer by stanbon(75887)

Which value of r gives the line passing through (3,2) and (r,-4) a slope of 3?
-------
Equation:
(-4-2)/(r-3) = 3
-----
r-3 = -6/3
r = -2+3
r = 1
------------
Cheers,
Stan H.
--------------